Solve the following system of linear equations: x + y + z = 2 (1) 6x − 4y + 5z = 31 (2) 5x + 2y + 2z = 13 (3) The solution of the system is the ordered triple: [1] (3, − 2, 1). [2] (3, 2, 1). [3] (0, 2, 0). [4] (1, 2, 1).


1
Expert's answer
2021-05-27T07:31:26-0400

"x + y + z = 2\ ......[1]
 \\ 6x - 4y + 5z = 31\ .....[2]
 \\ 5x + 2y + 2z = 13\ \ .....[3]"


By Substitution method:


Solve equation [1] for the variable  z 

"z = -x - y + 2"


Plug this in for variable  z  in equation [2]

"6x - 4y + 5\cdot(-x -y +2) = 31\\
 \Rightarrow x - 9y = 21"


Plug this in for variable  z  in equation [3]

"5x + 2y + 2\cdot(-x -y +2) = 13\\
 \Rightarrow 3x = 9"


 Solve equation [3] for the variable  x 

"3x = 9 
\\
 \Rightarrow \boxed {x = 3}"


 Plug this in for variable  x  in equation [2]

"(3) - 9y = 21\\
 \Rightarrow - 9y = 18"


Solve equation [2] for the variable  y 

"9y = - 18 \\

 \Rightarrow \boxed{ y = - 2 }"


 By now we know this much :

"x = 3\\
 y = -2\\
 z = -x-y+2"


Hence, Use the  x  and  y  values to solve for  z 

"z = -(3)-(-2)+2 = 1 \\\Rightarrow\boxed{z=1}"


Answer: [1] "\boxed{(x,y,z)=(3,-2,1)}"